# Pitfalls of Evidence-Based AI Policy

### v3

@@ 4.2 Ample Room for Progress
  • Brazil passed Bill No. 2338 of 2023 (Brazil, 2023) (enacted) on regulating the use of Artificial Intelligence, including algorithm design and technical standards in December 2024.
  • Canada recently established an AI Safety Institute (exists), and its proposed AI and Data Act (Canada, 2022) (proposed) is currently under consideration in House of Commons Committee.
@@
  • South Korea passed the [Act on the Development of Artificial Intelligence and Establishment of Trust (AI Basic Act) (National Assembly of the Republic of Korea, 2025) in December 2024.
@@
Brazil Canada China EU Korea UK USA 1. AI governance institutes O O* 2. Model registration O* 3. Model specification and basic info O O* O O 4. Internal risk assessments O* O 5. Independent 3rd-party risk assessments O O O 6. Plans to minimize risks to society O O O 7. Post-deployment monitoring reports 8. Security measures O 9. Compute usage O O* 10. Shutdown procedures O* O 11. Documentation availability O* O O O 12. Documentation comparison in court 13. Labeling AI-generated content O O 14. Whistleblower protections 15. Incident reporting
@@ 5 Building a Healthier Ecosystem
We often hear discussions about how policymakers need help from AI researchers to design technically sound policies. This is essential. But there is a two-way street. Policymakers can do a great deal to help researchers, governments, and society at large to better understand and react to AI risks.

### v4

@@ 4.2 Ample Room for Progress
  • Brazil proposed Bill No. 2338 of 2023 (Brazil, 2023) (proposed) on regulating the use of Artificial Intelligence, including algorithm design and technical standards. It passed the Brazilian senate in December 2024 but is pending further action before becoming law.
  • Canada recently established an AI Safety Institute (exists), and its AI and Data Act (Canada, 2022) (failed) recently failed in House of Commons Committee.
@@
  • South Korea passed the Act on the Development of Artificial Intelligence and Establishment of Trust (AI Basic Act) (National Assembly of the Republic of Korea, 2025) in December 2024.
@@
Brazil Canada China EU Korea UK USA 1. AI governance institutes O* O* 2. Model registration O* 3. Model specification and basic info O* O O 4. Internal risk assessments * O 5. Independent 3rd-party risk assessments O O O 6. Plans to minimize risks to society O* O O 7. Post-deployment monitoring reports 8. Security measures O 9. Compute usage O O* 10. Shutdown procedures * O 11. Documentation availability O O O 12. Documentation comparison in court 13. Labeling AI-generated content O O 14. Whistleblower protections 15. Incident reporting *
@@ 5 Building a Healthier Ecosystem
We often hear discussions about how policymakers need help from AI researchers to design technically sound policies. This is essential. But there is a two-way street. Policymakers can do a great deal to help researchers, governments, and society at large to better understand and react to AI risks (Kolt et al., 2024).